OpenStudy (anonymous):

if you have a sample of U- 235 with a mass of 50g and it has undergone 4 half-lives. how much did the original sample weigh ? please show step by step and the equation thanks ...

OpenStudy (anonymous):

After undergoing one **half life, how much mass is there if the original sample had mass M?

OpenStudy (anonymous):

if you have a sample of U- 235 with a mass of 50g and it has undergone 4 half-lives. how much did the original sample weigh .. this is the only information the question is asking

OpenStudy (anonymous):

Right, but do you know what "half life" means?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

can u help me solved this please i need to finished for homework

OpenStudy (anonymous):

^_^ If something has undergone 1 half life, then it had 1/2 (half) its original mass. \[m_1=\frac{1}{2}M\] If something has undergone 2 half lives, then it has half of half (1/2)(1/2) its original mass \[ m_2 = \frac{1}{2} m_1 = \frac{1}{2} \frac{1}{2} M = \frac{1}{4}M\] So what would it look like if it has undergone 4 half lives?

OpenStudy (anonymous):

sorry i dont get it at all

OpenStudy (anonymous):

my teacher give me a equation Mo(1/2)^n=MR

OpenStudy (anonymous):

mo= original mass ..n-#of half lives passed mr-remaining mass

OpenStudy (anonymous):

Right, exactly

OpenStudy (anonymous):

so if MR = 50g, and n=4, solve for Mo

OpenStudy (anonymous):

so x=(1/2)^n=50 g ?

OpenStudy (anonymous):

\[ \frac{M_0}{16} = M_R\] \[M_0 = 16 M_R\]

OpenStudy (anonymous):

\[ M_0 = 16*50 g\]

OpenStudy (anonymous):

i need an equation please

OpenStudy (anonymous):

\[ M_0\Big( \frac{1}{2} \Big) ^4 = M_R\] \[M_0 \frac{1}{16} = M_R\] \[M_0 = 16 M_R\] \[M_0 = 16 ( 50g)\] \[M_0 = 800 g\]

OpenStudy (anonymous):

thanks can u help me with anohter one

OpenStudy (anonymous):

Sure - the question asks for the weight of the sample though, and weight is mass times gravity.

OpenStudy (anonymous):

Thorium-234 has a half life of 24 days. If 360 days pass and you have only 0.5g of the originalsample, what was the mass of the original sample ?

OpenStudy (anonymous):

can u identify MR , N and mo for me please

OpenStudy (anonymous):

So what's MR in the question? (there's only one mass)

OpenStudy (anonymous):

is ask for the mass of the original sample

OpenStudy (anonymous):

Right, I get that.

OpenStudy (anonymous):

I'm asking *you

OpenStudy (anonymous):

y it say original is 0.5 bit it ask for original sample

OpenStudy (anonymous):

this is y im so confusing

OpenStudy (anonymous):

MR = 0.5 g It's 0.5 g after 360 days Half life is 24 days. 360 / 24 = ?

OpenStudy (anonymous):

360/24=15

OpenStudy (anonymous):

so N=15

OpenStudy (anonymous):

MR = 0.5g N = 15

OpenStudy (anonymous):

so is x(1/2)=0.5

OpenStudy (anonymous):

M0 (1/2)^15 = MR M0 (1/2)^15 = 0.5g M0 = 2^15 (0.5g)

OpenStudy (anonymous):

my teacher give usthe answer is 2x 10 ^4

OpenStudy (anonymous):

the the final answer is not correct compare to her did we do something wrong ?

OpenStudy (anonymous):

she gimy teacher giveme the answer but no work so i dont know how she get that

OpenStudy (anonymous):

2^15 = 32768 32768 * (.5g) = 16384g You only have 1 significant figure in the problem so 16384g rounds up to 20000g 20000g = 2 x 10^4g

OpenStudy (anonymous):

where u get 2?

OpenStudy (anonymous):

i thought .5^15 where u get 2 ^15 ?

OpenStudy (anonymous):

\[ M_0 \Big(\frac{1}{2} \Big)^{15} = M_R\] \[ M_0 \frac{1}{2^{15}} = M_R\] \[ M_0 = 2^{15}M_R\]

OpenStudy (anonymous):

yea but when i put in the equation how i plot it in ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

mo=(2)^15=0.5 ?

OpenStudy (anonymous):

\[M_0 = 2^{15}(0.5g)\]

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

r u a teacher u are good

OpenStudy (anonymous):

Just a tutor, but thanks! ^_^

OpenStudy (anonymous):

how can i contact u next time if i need help

OpenStudy (anonymous):

You can message, fan, or just tag me in your next question - hopefully the site will be finished with lots of its updates soon ^_^